# Line Cover and Related Problems

### v1

@@ 4 W[2]-hardness of covering points with subspaces
Consider an instance $(G=(V,E),k^{\prime})$ of Dominating Set. We assume without loss of generality that $V=[d]$, that $d,k^{\prime}>1$, and that no vertex in $G$ has degree d1d-1 (in which case the problem becomes trivial). For each $v\in[d]$, we create $dk^{\prime}$ points $\{\mathbf{p_{v,i}}\}_{i\in[dk^{\prime}]}$, each in $\mathbb{R}^{d}$, as follows: for each $i\in[dk^{\prime}]$, we set the $j^{th}$ coordinate of $\mathbf{p_{v,i}}$ as 1 (that is, $p_{v,i}[j]=1$) if $j\in N[v]$; the remaining entries will be set later. There are $d^{2}k^{\prime}$ points in total and we interpret them as an $d^{2}k^{\prime}\times d$ matrix $M$. Consider the $d^{2}k^{\prime}\times d^{2}k^{\prime}$ Vandermonde matrix $\mathbf{V}$ defined by the numbers $2,\dots,d^{2}k^{\prime}+1$. A key component of the reduction is the fact that the determinant of any square sub-matrix of $\mathbf{V}$ is non-zero [orucc2000explicit]. Consider columns $2$ to $d+1$ in $\mathbf{V}$, that is, the sub-matrix of $\mathbf{V}$ containing the columns from $2$ to 2+d12+d-1 and all rows. Now, we set any unset entries $M[i,j]$ as $\mathbf{V}[i,1+j]$. The matrix $M$ now defines the $d^{2}k^{\prime}$ points $\mathbf{p_{1,1}},\dots,\mathbf{p_{d,dk^{\prime}}}$ in the instance. Next, we set r=d1r=d-1, $B=0$, and $k=k^{\prime}$, that is, the number of allowed hyperplanes in a solution equals the number of allowed vertices in a dominating set. The constructed instance of Hyperplane Cover is therefore $(\{\mathbf{p_{1,1}},\dots,\mathbf{p_{d,dk}}\},k)$. Notice that even if we change the norm considered over the distances between points and their closest subspaces, the goal remains to find $k$ affine subspaces of dimension d1d-1 that cover all the points since the distances are nonnegative. We show the correctness of the reduction in the following lemma.
Lemma 1.
The instance $(G=(V,E),k^{\prime})$ is a yes-instance of Dominating Set if and only if $(\{\mathbf{p_{1,1}},\dots,\mathbf{p_{d,dk}}\},k)$ is a yes-instance of Hyperplane Cover.
@@
The forward direction follows directly from the construction. Let $S$ be a dominating set of size $k=k^{\prime}$. Then, the $k$ subspaces, each of dimension d1d-1, is given by $\{\{\mathbf{x}\mid x[i]=1\}\mid i\in S\}$. Since for any $v\in[d]$, there exists $i\in S$ such that $v\in N[i]$, we have that the subspace $x[i]=1$ covers all the points corresponding to $v$ (since they contain $1$ in the $i^{th}$ coordinate by construction).
For the reverse direction, let $S_{1},\dots,S_{k}$ denote the $k$ subspaces of dimension d1d-1 that cover the points, where the subspace $S_{j}$ is the set of points $\mathbf{x}=(x[1],\dots,x[d])$ that satisfy the equation $c_{j,0}+c_{j,1}x[1]+c_{j,2}x[2]+\dots+c_{j,d}x[d]=0$. Notice that we have a correspondence between the coefficients and the vertices in $G$: for each vertex $v\in V$, $c_{j,v}$ is the coefficient corresponding to $v$ in the linear equation defining $S_{j}$.
Let $(C_{1},\dots,C_{k})$ be an arbitrary $k$-partition of the $d^{2}k$ input points such that $C_{j}$ contains points that are covered by $S_{j}$. In the following claim, we show that we can assume more about the partition.
Claim 1.
Without loss of generality, we can assume the following for each vertex $v$ in the graph. All the $dk$ points corresponding to $v$, that is $\{\mathbf{p_{v,i}}\}_{i\in[dk]}$, belong to the same part: that is, there exists $j\in[k]$ such that $\mathbf{p_{v,i}}\in C_{j}$ for all $i\in[dk]$. Moreover, the non-zero coefficients in the linear equation defining $S_{j}$ is a non-empty subset of $N[v]\cup\{0\}$.
Proof.
Suppose that the partition $(C_{1},\dots,C_{k})$ does not satisfy the claim statement. By the pigeonhole principle, there is a part $C_{j}$ that contains $d$ points corresponding to $v$, say $\mathbf{p_{v,1}},\dots,\mathbf{p_{v,d}}$. Recall that the subspace $S_{j}$ is the set of points $\mathbf{x}=(x[1],\dots,x[d])$ that satisfy the equation $c_{j,0}+c_{j,1}x[1]+c_{j,2}x[2]+\dots+c_{j,d}x[d]=0$. The coefficients $c_{j,0},\dots,c_{j,d}$ satisfy the following system of $d$ linear equations.
$\begin{cases}c_{j,0}+c_{j,1}p_{v,1}[1]+c_{j,2}p_{v,1}[2]+\dots+c_{j,d}p_{v,1}[d]=0\\ \vdots\\ c_{j,0}+c_{j,1}p_{v,d}[1]+c_{j,2}p_{v,d}[2]+\dots+c_{j,d}p_{v,d}[d]=0.\end{cases}$ (1)
Let $(\textsc{NZ},\textsc{Z})$ be a partition of $\{0\}\cup[d]$ such that $q\in Z$ if and only if $c_{j,q}=0$. Equivalently, $(\textsc{NZ},\textsc{Z})$ partitions the coefficients depending on whether they are non-zero or zero. Firstly, we claim that $\textsc{NZ}\subseteq N[v]\cup\{0\}$. Suppose not, then since $\textsc{NZ}\neq\{0\}$ by the definition of an d1d-1 dimensional subspace, we infer that the set $\textsc{NZ}\setminus(N[v]\cup\{0\})$ is non-empty: let $q_{1},\dots,q_{\ell}\in[d]$ be the elements in $\textsc{NZ}\setminus(N[v]\cup\{0\})$. Note that $\ell<d$. Let $\tilde{c_{j}}=\sum_{q\in\textsc{NZ}\cap N[v]}c_{j,q}p_{v,1}[q]$. Then, the points $\{\mathbf{p_{v,1}},\dots,\mathbf{p_{v,\ell}},\mathbf{p_{v,\ell+1}}\}$ satisfy the following system of linear equations.
$\begin{cases}\tilde{c_{j}}+c_{j,q_{1}}p_{v,1}[q_{1}]+c_{j,q_{2}}p_{v,1}[q_{2}]+\dots+c_{j,q_{\ell}}p_{v,1}[q_{\ell}]=0\\ \vdots\\ \tilde{c_{j}}+c_{j,q_{1}}p_{v,\ell+1}[q_{1}]+c_{j,q_{2}}p_{v,\ell+1}[q_{2}]+\dots+c_{j,q_{\ell}}p_{v,\ell+1}[q_{\ell}]=0.\end{cases}$ (2)
The above system can also be expressed in the matrix form as follows.
$\begin{bmatrix}1&p_{v,1}[q_{1}]&\cdots&p_{v,1}[q_{\ell}]\\ 1&p_{v,2}[q_{1}]&\cdots&p_{v,2}[q_{\ell}]\\ \vdots&\vdots&\ddots&\vdots\\ 1&p_{v,\ell+1}[q_{1}]&\cdots&p_{v,\ell+1}[q_{\ell}]\\ \end{bmatrix}\begin{bmatrix}\tilde{c_{j}}\\ c_{j,q_{1}}\\ \vdots\\ c_{j,q_{\ell}}\end{bmatrix}=\begin{bmatrix}0\\ 0\\ \vdots\\ 0\end{bmatrix}$ (3)
Recall that the columns $q_{1},\dots,q_{\ell}$ correspond to vertices that are not adjacent to $v$. Thus, by construction, the first matrix, which forms a square sub-matrix of the $d^{2}k\times d^{2}k$ Vandermonde matrix $V$, has a non-zero determinant. Consequently, the above system is valid only when all of $\tilde{c_{j}},c_{j,q_{1}},\dots,c_{j,{q_{\ell}}}$ are zero. This is a contradiction to the fact that the coefficients are non-zero. Thus, we conclude that $\textsc{NZ}\subseteq N[v]\cup\{0\}$.
Since $\textsc{NZ}\subseteq N[v]\cup\{0\}$, we infer that all the points $\{\mathbf{p_{v,i}}\}_{i\in dk}$ are contained in $S_{j}$ (recall that by construction, all these points have the same value in the coordinates corresponding to $N[v]$). Thus, we can obtain a new $k$-partition from $(C_{1},\dots,C_{k})$ by moving all the points $\{\mathbf{p_{v,i}}\}_{i\in dk}$ to $C_{j}$. By repeatedly applying the above argument to each vertex in the graph, we obtain a partition satisfies the claim statement. ∎
Consider a $k$-partition $(C_{1},\dots,C_{k})$ that satisfies the statement in 1. Now we are ready to construct a dominating set $S^{*}$ in $G$. For each part $C_{j}$, we do the following. Let NZ denote the non-zero coefficients in the equation defining $S_{j}$. For each vertex $v$ such that the points corresponding to $v$ are contained in $C_{j}$, we know that $\textsc{NZ}\subseteq N[v]\cup\{0\}$. Consequently, we have that
$\textsc{NZ}\subseteq\bigcap_{v~:~\text{$\{\mathbf{p_{v,i}}\}_{i\in[dk]}\in C_{j}$}}\left(N[v]\cup\{0\}\right).$
Since $\textsc{NZ}\neq\{0\}$ by the definition of a (d1)(d-1)-dimensional subspace, we have that there is a non-zero index in the intersection of closed neighborhoods of all vertices $v$, whose points are contained in $C_{j}$. This implies that there is a vertex $w\in\textsc{NZ}\setminus\{0\}$ that is in the intersection of the closed neighborhoods of all vertices $v$ whose points are contained in $C_{j}$. We add $w$ to $S^{*}$. Since we add one vertex for each of the $k$ parts, we obtain a dominating set of size at most $k^{\prime}=k$ in $G$. ∎
Observe that the above reduction also rules out a $n^{o(k)}$-time algorithm for Hyperplane Cover. This is because such an algorithm would imply a $(d^{2}k^{\prime})^{o(k^{\prime})}=d^{o(k^{\prime})}$ algorithm for Dominating Set (on a $d$ vertex graph with solution size at most $k^{\prime}$) since we can always assume that $k^{\prime}\leq d$. This would refute the ETH [Chen20061346].
@@ 5 XP algorithm for Projective Clustering
If $R(\sigma)$ is not empty, then each of its non-empty semi-algebraically connected (which is equivalent to just connected on semi-algebraic sets as proven in [10.5555/1197095, Theorem 5.22]) components is a cell of $\mathcal{P}$ over $V$. Figure 6 shows an example of a family of polynomials on a single variable partitioning the $x$-axis into cells. For an algebraic set $W$, its real dimension is the maximal integer $d^{\prime}$ such that there is a homeomorphism of $0,1]^{d^{\prime}}$ in $W$. Naturally if $W\subset\mathbb{R}^{d}$, then $d^{\prime}<d$. The following theorem by Basu et al. [10.5555/1197095] gives an algorithm to compute a point in each cell of $\mathcal{P}$ over $V$.
Proposition 2 ([10.5555/1197095, Theorem 13.22]).
@@
The proof of Proposition 3 follows from the well-known reduction that allows reducing the problem of finding the affine subspace that minimizes the sum of squares of distances to a given point set to the problem of finding a subspace with the same property for a similar point set. Indeed, any $r$-dimensional affine space that minimizes the sum of squared perpendicular distances to the data points must pass through the centroid of these points (see, e.g., [Blum_Hopcroft_Kannan_2020]). Thus, by centering the data—i.e., by subtracting the centroid from each data point—we ensure that the best-fit affine subspace passes through the origin. Consequently, the problem of finding the best-fit $r$-dimensional subspace is identical to the classic principal component analysis (PCA). By the Eckart-Young theorem [Eckart1936TheAO], PCA can be solved efficiently via Singular Value Decomposition (SVD), which can be implemented in $\mathcal{O}(dn^{2}+d^{3})$ time [trefethen1997numerical].
@@
We distinguish between two cases: $r=0$ and $r>0$. In the first case, note that $n^{\mathcal{O}(dk(r+1))}=n^{\mathcal{O}(dk)}$. Moreover, the case is equivalent to $k$-Means. We can therefore use the $n^{\mathcal{O}(dk)}$-time algorithm of Inaba et al. [inaba1994applications].
@@ 6 Conclusion
In this work, we showed that Line Clustering is W[1]-hard when parameterized by $k$ and that Projective Clustering is in XP when parameterized by both $k$ and $d$. Complementing this result, we also show that the special case Hyperplane Cover is W[2]-hard when parameterized by $k$ alone (and it is known to be NP-hard for $d=2$).
We conclude with several open questions. Our first question concerns the $\operatorname{FPT}$-approximability of Projective Clustering. To the best of our knowledge, the existence of approximation algorithms with approximation ratio $f(d,k)$ and running time $g(d,k)\cdot n^{\mathcal{O}(1)}$ is open for any computable functions $f$ and $g$ only depending on $k$ and $d$. The parameterization by $k$ and $d$ also remains open. For the case when $r$ and $k$ are constants, Deshpande et al. [DBLP:journals/toc/DeshpandeRVW06] gave a polynomial-time approximation scheme (PTAS) for Projective Clustering. A rich body of literature exists on constructing coresets for specific cases of Projective Clustering (e.g.[DBLP:conf/stoc/FeldmanL11, DBLP:conf/fsttcs/VaradarajanX12]). For instance, when the input points have integer coordinates, small coresets can be constructed [DBLP:conf/fsttcs/VaradarajanX12], enabling randomized approximation algorithms that run in $f(k)\cdot n^{\mathcal{O}(1)}$ time. By combining our $\operatorname{XP}$ algorithm with such coreset constructions, efficient approximation algorithms can be developed for these cases. However, it is known that no $o(n)$-sized coreset exists for Projective Clustering, even in the special case where $r=k=2$ and d=3{d=3} [EV05, DBLP:conf/fsttcs/VaradarajanX12].
Another interesting open question concerns the existence of an exact $\operatorname{XP}$ algorithm for Projective Clustering parameterized by $d$ and $k$ for $\ell_{p}$-norms with $p\neq 2$. Already the base case with a single subspace differs with the case of $p=2$. Clarkson and Woodruff [ClarksonW15] show that for every $p\in[1,2)$, the problem of finding an $r$-dimensional subspace $F$ that minimizes (or even $(1+1/d^{\mathcal{O}(1)})$-approximates) $\sum_{i=1}^{n}\text{dist}(\mathbf{x}_{i},F)^{p}$ is NP-hard. For $p>2$, Deshpande et al. [DBLP:conf/soda/DeshpandeTV11] show NP-hardness and UGC-hardness.

### v2

@@ 4 $\operatorname{NP}$-hardness of covering points with two subspaces
Let $(W=\{w_{1},\dots,w_{2n}\},t)$ denote an instance of ECNP-d. We construct a point set $P$ in $\mathbb{R}^{n-1}$ as follows. For each $w_{i}\in W$, we add the point:
$\mathbf{p_{i}}=(v_{i},v_{i}^{2},\dots,v_{i}^{n-2},v_{i}^{n}),$
where $v_{i}=w_{i}-t/n$. Let $V=\{v_{1},\dots,v_{2n}\}$ denote the set of shifted values. Observe that an equal-cardinality partition in $W$ yields an equal-cardinality partition in $V$ where each part has sum 0. The total number of points in $P$ is $2n$. We claim that $P$ can be covered by exactly $2$ hyperplanes if and only if $V$ admits a valid equal-cardinality partition with sum 0 (which is true if and only if $W$ admits a valid equal-cardinality partition).
Suppose there exists a valid partition $(R_{1},R_{2})$ of $V$ where $|R_{1}|=|R_{2}|=n$ and the elements of each subset sum to exactly 0. We define two polynomials of degree $n$:
$Q_{1}(y)=\prod_{v\in R_{1}}(y-v)=y^{n}-\left(\sum_{v\in R_{1}}v\right)y^{n-1}+\dots+(-1)^{n}\prod_{v\in R_{1}}v$
and
$Q_{2}(y)=\prod_{v\in R_{2}}(y-v)=y^{n}-\left(\sum_{v\in R_{2}}v\right)y^{n-1}+\dots+(-1)^{n}\prod_{v\in R_{2}}v.$
Let the coefficients of $Q_{1}(y)$ be $a_{n},a_{n-1},\dots,a_{0}$ such that $Q_{1}(y)=\sum_{j=0}^{n}a_{j}y^{j}$, and let the coefficients of $Q_{2}(y)$ be $b_{n},b_{n-1},\dots,b_{0}$. Note that $a_{n-1}=b_{n-1}=0$ since the parts sum to 0. We define the hyperplane $H_{1}$ by the equation $a_{0}+a_{1}x_{1}+\dots+a_{n-2}x_{n-2}+a_{n}x_{n-1}=0$. Analogously, we define $H_{2}$ using the coefficients of $Q_{2}(y)$. By construction, evaluating $H_{1}$ on point $\mathbf{p_{i}}$ perfectly mirrors evaluating $Q_{1}(v_{i})$. Since the roots of $Q_{1}$ are exactly the elements of $R_{1}$, $H_{1}$ evaluates to 0 and safely covers $\mathbf{p_{i}}$ for all $v_{i}\in R_{1}$. Analogously, $H_{2}$ covers $\mathbf{p_{i}}$ for all $v_{i}\in R_{2}$. Thus, all $2n$ points of the form $\mathbf{p_{i}}$ are covered. The point set $P$ is entirely covered by $H_{1}$ and $H_{2}$.
Suppose $P$ can be covered by exactly two hyperplanes (in $\mathbb{R}^{n-1}$), $H_{1}$ and $H_{2}$. Let the equation for $H_{1}$ be $a_{0}+a_{1}x_{1}+\dots+a_{n-2}x_{n-2}+a_{n}x_{n-1}=0$ and $H_{2}$ be $b_{0}+b_{1}x_{1}+\dots+b_{n-2}x_{n-2}+b_{n}x_{n-1}=0$. We deliberately use $a_{n}$ and $b_{n}$ for the sake of presentation. For any point $\mathbf{p_{i}}$ covered by $H_{1}$, we have $a_{0}+a_{1}v_{i}+a_{2}v_{i}^{2}+\dots+a_{n-2}v_{i}^{n-2}+a_{n}v_{i}^{n}=0$. Thus, $v_{i}$ is a root of the polynomial
$Q_{1}(y)=a_{0}+a_{1}y+a_{2}y^{2}+\dots+a_{n-2}y^{n-2}+0\cdot y^{n-1}+a_{n}y^{n}.$
Since $H_{1}$ is a valid hyperplane, $(a_{0},\ldots,a_{n-2},a_{n})$ cannot be the zero vector; thus $Q_{1}(y)$ is not the zero polynomial. Since $Q_{1}(y)$ has degree at most $n$, and the numbers $v_{i}$ are distinct, it can have at most $n$ distinct roots. Consequently, $H_{1}$ can cover at most $n$ points of the form $\mathbf{p_{i}}$. By symmetric arguments, $H_{2}$ can also cover at most $n$ points of the form $\mathbf{p_{i}}$. We define the polynomial $Q_{2}(y)$ analogously. Since there are $2n$ points of the form $\mathbf{p_{i}}$ in total, $H_{1}$ is forced to cover exactly $n$ such points (corresponding to subset $R_{1}$), and $H_{2}$ must cover exactly the remaining $n$ such points (subset $R_{2}$). This guarantees that $a_{n}\neq 0$ and $b_{n}\neq 0$ (otherwise the degree drops and they could not cover $n$ points). The sum of the roots of $Q_{1}(y)$ is given by: $a_{n-1}/a_{n}=0$. Thus, we infer that $\sum_{v\in R_{1}}v=0$. Consequently, the corresponding original elements $w_{i}=v_{i}+t/n$ sum to $\sum_{v_{i}\in R_{1}}(v_{i}+t/n)=0+n(t/n)=t$, and we obtain the partition that forms a valid solution to the ECNP-d instance. ∎
@@ 5 XP algorithm for Projective Clustering
If $R(\sigma)$ is not empty, then each of its non-empty semi-algebraically connected (which is equivalent to just connected on semi-algebraic sets as proven in [5, Theorem 5.22]) components is a cell of $\mathcal{P}$ over $V$. Figure 6 shows an example of a family of polynomials on a single variable partitioning the $x$-axis into cells. For an algebraic set $W$, its real dimension is the maximal integer $d^{\prime}$ such that there is a homeomorphism of $0,1]^{d^{\prime}}$ in $W$. Naturally if $W\subset\mathbb{R}^{d}$, then $d^{\prime}<d$. The following theorem by Basu et al. [5] gives an algorithm to compute a point in each cell of $\mathcal{P}$ over $V$.
Proposition 1 ([5, Theorem 13.22]).
@@
The proof of Proposition 2 follows from the well-known reduction that allows reducing the problem of finding the affine subspace that minimizes the sum of squares of distances to a given point set to the problem of finding a subspace with the same property for a similar point set. Indeed, any $r$-dimensional affine space that minimizes the sum of squared perpendicular distances to the data points must pass through the centroid of these points (see, e.g., [6]). Thus, by centering the data—i.e., by subtracting the centroid from each data point—we ensure that the best-fit affine subspace passes through the origin. Consequently, the problem of finding the best-fit $r$-dimensional subspace is identical to the classic principal component analysis (PCA). By the Eckart-Young theorem [21], PCA can be solved efficiently via Singular Value Decomposition (SVD), which can be implemented in $\mathcal{O}(dn^{2}+d^{3})$ time [45].
@@
We distinguish between two cases: $r=0$ and $r>0$. In the first case, note that $n^{\mathcal{O}(dk(r+1))}=n^{\mathcal{O}(dk)}$. Moreover, the case is equivalent to $k$-Means. We can therefore use the $n^{\mathcal{O}(dk)}$-time algorithm of Inaba et al. [33].
@@ 6 Conclusion
In this work, we showed that Line Clustering is $\operatorname{W}[1]$-hard when parameterized by $k$ and that Projective Clustering is in $\operatorname{XP}$ when parameterized by both $k$ and $d$. Complementing this result, we also show that the special case Hyperplane Cover is $\operatorname{NP}$-hard even when $k=2$ (and it is known to be $\operatorname{NP}$-hard for $d=2$).
We conclude with several open questions. Our first question concerns the $\operatorname{FPT}$-approximability of Projective Clustering. To the best of our knowledge, the existence of approximation algorithms with approximation ratio $f(d,k)$ and running time $g(d,k)\cdot n^{\mathcal{O}(1)}$ is open for any computable functions $f$ and $g$ only depending on $k$ and $d$. The parameterization by $k$ and $d$ also remains open. For the case when $r$ and $k$ are constants, Deshpande et al. [17] gave a polynomial-time approximation scheme (PTAS) for Projective Clustering. A rich body of literature exists on constructing coresets for specific cases of Projective Clustering (e.g.[24, 46]). For instance, when the input points have integer coordinates, small coresets can be constructed [46], enabling randomized approximation algorithms that run in $f(k)\cdot n^{\mathcal{O}(1)}$ time. By combining our $\operatorname{XP}$ algorithm with such coreset constructions, efficient approximation algorithms can be developed for these cases. However, it is known that no $o(n)$-sized coreset exists for Projective Clustering, even in the special case where $r=k=2$ and d=3{d=3} [22, 46].
Another interesting open question concerns the existence of an exact $\operatorname{XP}$ algorithm for Projective Clustering parameterized by $d$ and $k$ for $\ell_{p}$-norms with $p\neq 2$. Already the base case with a single subspace differs with the case of $p=2$. Clarkson and Woodruff [14] show that for every $p\in[1,2)$, the problem of finding an $r$-dimensional subspace $F$ that minimizes (or even $(1+1/d^{\mathcal{O}(1)})$-approximates) $\sum_{i=1}^{n}\text{dist}(\mathbf{x}_{i},F)^{p}$ is $\operatorname{NP}$-hard. For $p>2$, Deshpande et al. [18] show $\operatorname{NP}$-hardness and UGC-hardness.